Both C and D PLEASE!!!Course Home <Gausss Law Exercise 22.17 Part C What is the magnitude of the net electric field at point y 0.620 m on the y-axis? A very long uniform line of charge has charge per unit length 4.76 |uC/m and lies along the x-axis. A second long uniform line of charge has charge per unit length-2.44 μC/m and is parallel to the x-axis at y0.400 m. E= N/C SubmitP Previous Answers Request Answer X Incorrect; Try Again; 5 attempts remaining The correct answer does not depend on: EO- Part D What is the direction of the net electric field at point y3 0.620 m on the y-axis?
Answer #1

c)

E = 1/ ( 2pi epsilon_not) * ( L1 / r1 + L2 / r2)

E = 1/ ( 2*3.14* 8.85*10^-12)* ( - 2.44 / (0.62 - 0.4) + 4.76 / 0.62)*10^-6

E = 6.14*10^4 N/C

=======

d)

directed along - y axis.

======

do comment in case any doubt, will reply for sure.. Goodluck